The sum of Nora's age and her grandmas age is 65. Four time law Nora's age is 10 less than her grandmas age. Fund Nora's present age and her grandmas present age.
Answer by stanbon(75887)

The sum of Nora's age and her grandmas age is 65. Four time law Nora's age is 10 less than her grandmas age. Fund Nora's present age and her grandmas present age.
------
Equations:
n + g = 65
4n = g-10
-----
Rearrange:
n + g = 65
4n- g =-10
----
Add and solve for "n":
5n = 55
Nora's age = 11 years
----
solve for "g":
n + g = 65
11 + g = 65
grandma's age is 54
==========================
Cheers,
Stan H.
==========================